a man uses a 900w electric iron to press his cloth for an average of 4 hours a week for 5 weeks . if the cost of electrical energy is 30 kobo per unit , calculate the cost of energy for pressing


1
Expert's answer
2021-06-17T15:09:31-0400

The total amount of energy can be calculated as the product of the power and the time the iron worked.

So the energy will be "900\,\mathrm{W}\cdot 4\,\mathrm{h}\cdot5 = 18000\,\mathrm{W\cdot h} = 18\,\mathrm{kW h}" .

Therefore, the cost will be "18\,\mathrm{kW h} \cdot 30\,\mathrm{k/kWh} = 540\,\mathrm{k}."